OpenStudy (anonymous):

Find the prime factorization of each number. If the number is prime, state this. 39, 30, 27, 150, 40, 31, 210, 33. If you can, explain the steps you did to at least of the above numbers. Thanks in advance :)

OpenStudy (anonymous):

39=3*13 30=2*3*5 27=3^3 150=2*3*5*5

OpenStudy (anonymous):

what does ^ mean?

OpenStudy (anonymous):

40=2*2*2*5 31 is prime 210=2*3*5*7 33=3*11

OpenStudy (anonymous):

^ means exponent

OpenStudy (anonymous):

27=3*3*3=3^3

OpenStudy (anonymous):

Okay, how did you find these? Can you teach me?

OpenStudy (anonymous):

For example, how would i find the prime factorization of 115?

OpenStudy (anonymous):

do u know what r the prime nos?

OpenStudy (anonymous):

can u list them?

OpenStudy (anonymous):

2, 3, 5, 7, 11, 13, 17, 19, 23 & so on..

OpenStudy (anonymous):

right :)

OpenStudy (anonymous):

Okay, but i dont understand what im suppose to do with them lol

OpenStudy (anonymous):

now for 115 u have to see by which prime no is 115 divisible?

OpenStudy (anonymous):

it ends at 5,so divisible by 5

OpenStudy (anonymous):

so do i start with 2 & go up?

OpenStudy (anonymous):

yeah, but 115 isnt divisible by 2

OpenStudy (anonymous):

115/5=23...ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

23 is prime

OpenStudy (anonymous):

now see 23 is prime, so we r end up 115=5*23

OpenStudy (anonymous):

Oh, okay.. i get it now.